# DeepSeek vs. ChatGPT: A Comparative Study for Scientific Computing and Scientific Machine Learning Tasks

### v1

@@ 3 Summary
In this study, we evaluated and compared the performance of recently released large language models from OpenAI (ChatGPT 4o, ChatGPT o3-mini-high) and DeepSeek (DeepSeek V3, DeepSeek R1) for diverse tasks in computational mathematics and scientific machine learning. We designed challenging problems that required advanced mathematical reasoning and domain-specific knowledge, including numerical integration, finite difference methods (FDM), finite element methods (FEM), and scientific machine learning tasks such as image recognition, physics-informed neural networks (PINNs), and Deep Operator Networks (DeepONet). Our focus was on assessing the models’ ability to select appropriate numerical methods or neural network architecture and implement them correctly in Python. The results show that reasoning-optimized models, DeepSeek R1 and ChatGPT o3-mini-high, consistently performed better in recognizing the nature of the problem and making their decisions accordingly. In fact, many of their choices are similar to what we would do in solving these problems both for scientific computing and scientific machine learning. In contrast, general-purpose models, DeepSeek V3 and ChatGPT 4o, sometimes fail to account for specific properties of the problem or instructions from the user, consequently generating incorrect or inaccurate solutions.
Our study highlights the growing usefulness of LLMs in scientific research, as well as the increasing competition between OpenAI and DeepSeek models, as both developers continue to refine their models for tasks in mathematics, coding, and scientific computing. Our findings also expose the limitations of these LLMs, as they may respond to user instructions with ambiguous or incorrect solutions, which could further confuse a human researcher. Our findings highlight the need for continued improvements in LLMs for scientific problem-solving. Future work should explore additional benchmarking methods and assess LLMs on more complex, real-world computational challenges, where a cascade of decisions is required at the different stages of a project. By doing so, researchers can make more informed decisions on when and how to use LLM assistants effectively and responsibly in their daily work.

### v2

@@ 3 Summary
In this study, we evaluated and compared the performance of recently released large language models from DeepSeek (DeepSeek V3, DeepSeek R1), OpenAI (ChatGPT 4o, ChatGPT o3-mini-high), and Anthropic (Claude 3.7 Sonnet, Claude 3.7 Sonnet with extended thinking) for diverse tasks in scientific computing and scientific machine learning. We designed challenging problems that require advanced mathematical reasoning and domain-specific knowledge, including numerical integration, finite difference methods (FDM), finite element methods (FEM), and scientific machine learning tasks such as image recognition, physics-informed neural networks (PINNs), and Deep Operator Networks (DeepONet). Our focus was on assessing the models’ ability to select appropriate numerical methods or neural network architecture and implement them correctly in Python. The results show that reasoning-optimized models, DeepSeek R1, ChatGPT o3-mini-high, and Claude with extended thinking, consistently performed better in recognizing the nature of the problem and making their decisions accordingly. In fact, many of their choices are similar to what we would do in solving these problems both for scientific computing and scientific machine learning. In contrast, general-purpose models, DeepSeek V3 and ChatGPT 4o, sometimes fail to account for specific properties of the problem (such as stiffness) or instructions from the user (such as to implement from scratch), consequently generating incorrect solutions.
Our study highlights the growing practicality of using LLMs in scientific research, as well as the increasing competition between DeepSeek, OpenAI, and Anthropic, as all front-running developers continue to refine their models for tasks in mathematics, coding, and scientific computing. Our findings also expose the limitations of these state-of-the-art LLMs, demonstrated by their ambiguous or incorrect responses, which could confuse a human researcher unfamiliar with the subject. Our findings highlight the need for continued improvements in LLMs for scientific problem-solving. Future work should explore additional benchmarking methods and assess LLMs on more complex, real-world computational challenges, where a cascade of decisions is required at the different stages of a project. By doing so, researchers can make more informed decisions on when and how to use LLM assistants effectively and responsibly in their daily work.